OpenStudy (anonymous):

Quadratic Factoring! 7x^2 + 40x - 25 = 0

OpenStudy (anonymous):

and 35x^2 - 69x + 28 =0

OpenStudy (anonymous):

first one does not factor using integers. use the quadratic formula to solve

OpenStudy (anonymous):

@satellite73 I'm not allowed to,our teacher wants us to turn it into something like (x+6)(7x+4)

OpenStudy (anonymous):

well you can't maybe there is a typo

OpenStudy (anonymous):

make sure the problem is exactly \[7x^2+40x-25=0\] if it is you can use the quadratic formula to find the zeros, but you will not be able to factor until you have them, because they are irrationals numbers

OpenStudy (anonymous):

Well what about this: 7x^2 + 32 = 7 - 40x and 42x^2 - 69x + 20 = 7x^2 - 8 Those are the very raw original problems beforeI put them over to one side to equal zero like my teacher wants us to do.

OpenStudy (anonymous):

\[7x^2 + 32 = 7 - 40x \iff7x^2+40x+25=0\]

OpenStudy (anonymous):

add \(40x\) subtract \(7\) now you can factor

OpenStudy (anonymous):

try \[(7x+5)(x+5)=0\]

OpenStudy (anonymous):

@satellite73 Ok :) I think that will work. Thank you!

OpenStudy (anonymous):

yw